OpenStudy (anonymous):

Kaye multiplies the numbers 4,5,6,7,8 and 9. Leon multiplies the numbers 5,6,7 and 8 together. Marie says she can divide Kaye's answer by Leon's answer in her head. Marie's answer should be:

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

36

OpenStudy (anonymous):

\[\frac{4.5.6.7.8.9}{5.6.7.8}=4*9=36\]

OpenStudy (anonymous):

ok thank you